OpenStudy (anonymous):

passing through (5,6) with an x intercept of -5 what is the point slope and slope intercept?

OpenStudy (anonymous):

OK. The x-intercept is whenever you let y=0. Imagine the x-intercept as the intersection of the x-axis. That means the y-values are always zero. That being said, you'll have another pair of coordinates (-5, 0). So find the slope of the points (5, 6) and (-5, 0) by using the slope formula which of course you should know as m=(y2-y1)/(x2-x1). So the slope is 3/5. The point slope formula is y-y1=m(x-x1). So pick any coordinates and plug them in, which I'll choose (-5, 0) because it's easier. So the answer in slope intercept form or y=mx+b is y=(3/5)x+(3/25). I hope this helps!!

OpenStudy (anonymous):

yeah it helps a lot thanks. I had half of this correct.